x^2-9x-36=0 solve for x
Answer by jessica43(140)

To solve this problem you need to factor the trinomial, which is the opposite of the FOIL method. The factors will have this form:
(x A)(x B)
where A and B will be the factors of the third term (-36). The factors of -36 are: -1, 1, -2,2,-3, 3,-4, 4,-6,6 -9, 9, -12, 12,-18, 18, -36 and -36. So these are our possibilities of A and B: -1&36, 1&-36, -2&18, 2&-18, -3&12, 3&-12, and so on.
Now, we need to see which of these combinations, when added, will equal -9 (since the second term of our trinomial is -9x).
The only combination that could equal -9 when added is -12 and 3.
So:
x^2 - 9x - 36 = 0
(x-12)(x+3)= 0
So for this to equal zero, x = 12 or x = -3.
Your answers are x = 12 and x = -3.